# Sorting inversion sequences

### v1

@@ 4.2 The generating tree for the pop stack sortable inversion sequences
As in other cases, one can show that the generating tree (based on Algorithm [16]) of $I_{n}(120,201,1010)$, that is, of the pop stack sortable inversion sequences, is given by a root $a_{1}$ and the rules
@@
For any sequence of nodes $f_{m,j}$, we define $F_{m,j}(x)$ as the generating functions for the number of nodes in the subtrees $\mathcal{T}(B;f_{m,j})$, where $f\in\{a,b,c,d\},F\in\{A,B,C\}$. Define
@@
$C(v,u)$ $=\frac{x}{(1-v)(1-uv)}+\frac{x}{vu(1-v)}(C(v,u)-C(v,0))$
$+\frac{x}{v^{2}u^{2}(1-v)}(A(uv)-A(0)-\frac{uv}{x}A(0)+vu+vuB(vu,1)).$
@@ 5.   Open Problems
There are many directions left to continue the study of sortable inversion sequences. The most obvious question is the elusive enumeration problem for the inversion sequences of length $n$ which avoid $120$. Already interesting as a standalone pattern avoidance question, the fact that these inversion sequences are exactly those sortable by the classic stack sorting algorithm by Knuth [15] increases the intrigue.
@@
How many inversion sequences of length $n$ are stack sortable, i.e. avoid $120$?
We have classified the pop stack sortable inversion sequences and words and provided a generating tree for the inversion sequence case in subsection 4.2. However, the combinatorial enumeration is still open.
@@
Another avenue of exploration is considering other sorting machines or algorithms. For instance, the authors considered a sorting algorithm on a single stack prioritizing outputting the next correct entry of the sequence until no other moves were available and then either returning the elements in order to be sorted again [22] or reversing the output for the next pass [23]. In particular, the $2$-reverse-pass sortable permutation class has a reasonably small basis, namely $\{2413,2431,23154\}=\{1302,1320,12043\}$ that may prove tractable.

### v5

@@ 4.2 The generating tree for the pop stack sortable inversion sequences
As in the previous case, one can show that the generating tree (based on the algorithm given in [20]) of $I_{n}(120,201,1010)$, that is, of the pop stack sortable inversion sequences, is given by a root $a_{1}$ and the rules
@@
For any sequence of nodes $f_{m,j}$, we define $F_{m,j}(x)$ as the generating functions for the number of nodes in the subtrees $\mathcal{T}(B;f_{m,j})$, where $f\in\{a,b,c\},F\in\{A,B,C\}$. Define
@@
$C(v,u)$ $=\frac{x}{(1-v)(1-uv)}+\frac{x}{uv(1-v)}(C(v,u)-C(v,0))$
$+\frac{x}{u^{2}v^{2}(1-v)}(A(uv)-A(0)-uv\frac{\partial}{\partial v}A(v)\mid_{v=0})+\frac{x}{uv(1-v)}(B(vu,1)-B(0,0)).$
@@
4.3 Inversion sequences sortable by a pop stack of depth two
Recall from Corollary 3.11 that the inversion sequences sorted by a pop stack of depth two are precisely those which avoid $P=\{120,201,210,1010\}$. We can define the generating tree $T_{P}$ for inversion sequences that avoid $P$ as having a root $a_{1}$ and satisfying the following rules:
$a_{m}$ $\rightsquigarrow a_{m+1},b_{m,1},\ldots,b_{m,m},$
$c_{m}$ $\rightsquigarrow c_{m+1},a_{m+2},b_{m+1,1},\ldots,b_{m+1,m+1},$
where $a_{m}=0^{m},b_{m,j}=0^{m}j$, and $c_{m}=0^{m}10$. We label the inversion sequence 00 by $a_{1}$, so the root of the tree $T_{P}$ is indeed $a_{1}$. Now, let us show that the succession rules hold in $T_{P}$. Based on the algorithm introduced in [17] (for more examples and proofs, see [7, 8]), we have the following:
  • The children of $a_{m}$ in $T_{P}$ are $a_{m}j$ with $j=0,1,\ldots,m$, which are equal to $a_{m+1},b_{m,1},\ldots,b_{m,m}$. Hence, the first succession rule holds.
  • The children of $b_{m,j}$ in $T_{P}$ are $0^{m}ji$ with $i=0,1,\ldots,m+1$. But, note that the set of inversion sequences of the form $0^{m}ji\pi\textquoteright$ ’ with $10\leq i\leq j-$ 1 that avoid $P$ can be mapped bijectively to the set of inversion sequences of the form $0^{m+1-j}10\pi^{\prime\prime}$ that avoid $P$ (by mapping each letter $i$ to 00 and each letter $s\geq j$ to $s+1-j$ and writing the first initial $0^{m}$ as $0^{m+1-j}$). Hence, $T_{P}(0^{m}ji)\cong T_{P}(c_{m+1+j})$, for all $i=0,1,\ldots,j-1$.
    Now, let us look at the remaining children of $b_{m,j}$ in $T_{P}$, which are $0^{m}ji$ with $i=j,j+1,j+2,\ldots,m+1$. Clearly, an inversion sequence $0^{m}jj\pi\textquoteright$ ’ avoids $P$ if and only if the inversion sequence $0^{m+1}j\pi^{\prime}$ avoids $P$, so $T_{P}(0^{m}jj)\cong T_{P}(0^{m+1}j)$. Next consider the inversion sequence $\pi=0^{m}ji\pi\textquoteright$ ’ with $j+1\leq i\leq m+1$. Notice that $\pi$ avoids $P$ if and only if $0^{m+1-j}(i-j)\pi^{\prime\prime}$ avoids $P$ (by mapping each letter $s\geq j$ to $s-j$ and writing the initial $0^{m}j$ as $0^{m+1-j}$). Hence, the children of $b_{m,j}$ in $T_{P}$ are $(c_{m+1-j})^{j},b_{m+1,j},b_{m+1-j,1},\ldots,b_{m+1-j,m+1-j}$, which proves the second succession rule.
  • The children of $c_{m}$ in $T_{P}$ are $0^{m}100,0^{m}101,\ldots,0^{m}10(m+2)$. Clearly, an inversion sequence $0^{m}100\pi\textquoteright$ ’ avoids $P$ if and only if $0^{m+1}10\pi\textquoteright$ ’ avoids $P$. Also, an inversion sequence $0^{m}10i\pi\textquoteright$ ’ avoids $P$ if and only if $0^{m+2}\pi^{\prime\prime}$ avoids $P$ (here $\pi^{\prime\prime}$ is obtained by mapping each letter $s$ in $i\pi\textquoteright$ ’ to $s-1$). Thus the children of $c_{m}$ in $T_{P}$ are $c_{m+1},a_{m+2},b_{m+1,1},\ldots,b_{m+1,m+1}$, which shows that the third succession rule holds.
@@ 5.   Open Problems
There are many directions left to continue the study of sortable inversion sequences. The most obvious question is the elusive enumeration problem for the inversion sequences of length $n$ which avoid $120$. Already interesting as a standalone pattern avoidance question, the fact that these inversion sequences are exactly those sortable by the classic stack sorting algorithm by Knuth [19] increases the intrigue. The first author and Shattuck [25, Section 4] obtained a recurrence relation for these inversion sequences yielding a functional equation in three variables (sequence number A263778 in OEIS [32]). We note that the construction method used here was later generalized by Testart [35] as part of his work to complete the enumeration of 22 open cases of pattern-avoiding inversion sequences.
@@
Is there a nicer combinatorial formula for the number of inversion sequences of length $n$ are stack sortable, i.e. avoid $120$?
Next, while we have classified the pop stack sortable inversion sequences and words and provided a generating tree for the inversion sequence case in Section 4.2, the following combinatorial enumeration is also still open.
@@
Another avenue of exploration is to consider other sorting machines or algorithms. For instance, the authors considered a sorting algorithm on a single stack prioritizing outputting the next correct entry of the sequence until no other moves were available and then either returning the elements in order to be sorted again [26] or reversing the output for the next pass [27]. In particular, the $2$-reverse-pass sortable permutation class has a reasonably small basis, namely $\{2413,2431,23154\}=\{1302,1320,12043\}$ that may prove tractable.